Calculate the molar mass of acetaminophen (C_{8}*H_{9} * N * O_{2}) an over-the-counter pain reliever also known as paracetamol.

1
Expert's answer
2022-02-10T12:53:03-0500

"C=12,H=1, N=14, O=16"

molar mass ="12\times 8+1\times 9+14+ 16\times 2=151g"